Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
vardaro, bonavena, delaurentis, zendri - 23:03 Friday 08 December 2023 (62684) Print this report
FIS Injection with SR Misaligned

This is a brief entry the full one will follow in the next days.

Today we injected SQZ in normal configuration: SR aligned ando also with SR Misaligned by +1.1urad in TY.

The SR misalignment was completed at 17:22 UTC. Then we started the characterization with Injections and a phase scan.

At 18:43 we started a long run of SQZ injected with SR misaligned

At 18:50 we put the ITF in LN3

At 19:00 UTC we put the ITF STATUS node in autorelock filesafe

The maximum range reached during the injection was 48.8 Mpc.


We left the system in the following way:

  • ITF_STATUS: Autorelock filesafe
  • SQB1_Shutter: Open
  • SR_TY: Misaligned by +1.1urad 
Images attached to this report
Comments to this report:
vardaro - 14:35 Saturday 09 December 2023 (62686) Print this report

I rotated the SQB1_HWP2 to send back SQZ to HD detector

I measured 5.85 dB of SQZ and 8.35 dB of ASQZ. With 8.8 dB of generated squeezing.

I asked to Henning to switch back in BAB. The SQB1 retroreflector is still on.

vardaro - 0:16 Monday 11 December 2023 (62695) Print this report

The retroreflector is removed.

I did manually FMODERR (sidebands are off)

Then I unclipped the IR in the faraday.

I left the system in FLT_IR_AA_FMODERR_TUNED

Images attached to this comment
bonavena, delaurent, vardaro, zendri - 17:30 Tuesday 12 December 2023 (62713) Print this report

The shift aims to inject the FIS into the ITF. The idea was to work in LOW_NOISE_2 because in this state we had all the LSC dither lines on and all the slow tuning loops of the ITF on. This ITF state was useful during the SR angular position movement. We switched to LOW_NOISE_3 only at the end of the shift.

Preparatory work

  • At the beginning of the shift we checked the SQZ level generated with the HD Detector on EQB1, i.e. we measured the EQB1_HD_DIFF_RF_4MHz_mag signal  @UTC 9:36.  (4 MHz_max = 8.34 mV, 4MHz_min = 3.14mV; i.e. 8.27 dB)

  • @ UTC 9:42 We rotated the HWP2 on SQB1 to send the SQZ beam toward the ITF.

  • @UTC 9:53 We closed the AA loop on B1 4 MHz.

  • @ UTC 10:00 The LO was shuttered and the final position of the slow shutter on SQB1 was -1600000 steps to maximize the B1 4 MHz mag.

Problem with the CC phase

When we closed the CC loop and the AA loop of SQZ into the ITF (on B1) we realized that the squeezing phase was unstable. Initially, we suspected a problem in the computation of the demodulation phase by the DAQ. Thus we contacted Alain and we did the following checks:

  • We looked at the SQZ_CC_8MHz_mag, dfreq and dphi and we did not notice any problem in the SQZ 8MHz reference readout

  • We checked the fluctuation of the CC phase into the ITF by looking the B1_PD1_4MHz_Q and mag and we realized a strong correlation with the SQZ_PumpPhase_Error channel

  • We cross-checked also with the EQB1_HD_Detector trying to measure the SQZ level with it but we found the same problem.

  • We realized that the pump phase Error signal was not more around zero and that the SQZ_PumpPhase_HV_Monitor channel was not feeding any voltage to the PZT HV amplifiers

  • We realized that the PumpPhase Realay on the rack was disconnected. We reconnected it and it started to work all normally

N.B. This issue was because we worked with SQZ_MAIN in READY_FOR_FIS_INJECTION. The right procedure is work with SQZ_MAIN in SQZ_LOCKED

After that, we could properly inject SQZ into the ITF.

FIS Injection with SR aligned

  • @ UTC 11:45 We took 5 min of shot noise reference.

  • @UTC 11:51 We set the SQZ_MAIN Metatron node in SQZ_LOCKED and we run a CC phase scan with the script /virgoDev/AEI_SQZ/  (strange not reliable image phase scan01). At the beginning of the scan the filter cavity unlocked and an huge glitch spoiled the sensitivity in the first part. Thus we repeated it.

  • @UTC 11:56 we performed another CC phase scan but @12:03 UTC the ITF unlocked during it

  • @UTC 13:35 we could perfom a good phase scan (Figure 2). The last part was spoiled by huge glitches. Probably the filter cavity unlocked again. During the scan we measured 4.2 dB of SQZ. We measured between 4.2dB and 5 dB of ASQZ (2.7rad) and between 0.8 dB and 1.1 dB of SQZ(1.3 rad) depending from the observed frequency band

  • During the phase scan we looked at the channel SQZ_B1_PD1_4MHz_I_rms that corresponds to the in-loop residual phase noise of the CC loop and we observed that its value changed by more than a factor 2 during the scan (as expected). We tried thus to optimize the gain. We decided to not perform a noise injection (next time could be useful) but we tried to test different gains. Starting from 14:13 UTC we explored a region between 15000 and 100000 and we found an optimal gain at 80000. We did not observe any difference in the Hrec BRMS and in the detected SQZ level. From now on we worked with CC gain =80000 and we saved it in the SQZ_MAIN.ini config file.

  • @UTC 14:44:30  we performed another phase scan with the CC GAIN = 80000 and CC dither line off we found a strange behaviour in the data probably due to a glitch and we dropped the data.

  • We tried to run another CC phase scan but @15:10 UTC the ITF unlocked

At 15:40 UTC the ITF relocked. We decided to not to perform other phase scan but to take SQZ, ASQZ and Shot noise reference, then misalign the SR mirror (figure 3)

  • @UTC 15:59 After 15 min the ITF locked and we started acquiring 2 minutes of shot noise

  • @UTC 16:06  Acquisition of 5 minutes in SQZ (1.1 rad) 

  • @UTC 16:13  Acquisition of 5 minutes in ASQZ (2.7 rad) 

  • @UTC 16:20  Acquisition stopped

FIS Injection with SR misaligned

  • @UTC 16:42:50 We started the SR mirror misalignment procedure: we open the SR angular controls and we moved it by steps of 0.1 urad with the GUI. 

  • @UTC 17:22 SR misalignment finished (11 steps) TY +1.1urad (Figure 4)

Once the SR was misaligned we used the same phase for SQZ, ASQZ that we used for SR aligned:

  • @UTC 17:30 ASQZ injection 2.7 rad( 5 min). We measured about 3dB of ASQZ

  • @UTC 17:37 SQZ 1.1 rad (5 min). We measured about 0 dB of SQZ

  • @UTC 17:56 we started a phase scan + shot noise reference (Figure 5).

After the scan we left the system with SQZ injected 

  • @UTC 18:43  SQZ injection  0.666 rad

  • @UTC 18:51  LOW NOISE 3 + SQZ

  • @UTC 19:00 ITF Status in autorelock failsafe 

  • @UTC 21:36 +0.1urad with SR TY

  • @UTC 21:40 0.8rad CC phase

  • @UTC 22:12 0.5rad CC phase

  • @UTC 22:40 SQZ Off (Figure 6)

  • @UTC 22:57 SQZ Injected no CC dither line

At the end of the shift, we realized that the glitches were due to the unlock of FLT and usually were caused by a mistuned threshold (LockFlag) in the automation. We fixed it we changed the maxGR level from 6V to 5.5V in the SQZ_FLT.ini file.

Images attached to this comment
Search Help